OpenStudy (anonymous):

Find a fourth-degree polynomial that has the indicated zeros: _ _ -2/5 (multiplicity 2), 3+/2, and 3-/2.

OpenStudy (anonymous):

\[(5x+2)^2\] is one factor. that makes \[-\frac{2}{5}\] a zero with multiplicity 2

OpenStudy (anonymous):

the other factor is \[(x-(3+\sqrt{2}))(x-(3-\sqrt{2}))\] so you have to multiply that out, which is easier than it looks. can you do it?

OpenStudy (anonymous):

\[x^2-6x+7\]

OpenStudy (anonymous):

?right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so now if you need to you have to multiply \[(x^2-6x+7)(5x+2)^2\] to put in standard form. if you need to